# CTIBench: A Benchmark for Evaluating LLMs in Cyber Threat Intelligence

### v1

@@ 1 Introduction
Addressing this gap, we propose CTIBench, a novel suite of benchmark tasks and datasets to evaluate LLMs in cyber threat intelligence. To this end, we first create a knowledge evaluation dataset (CTI-MCQ) with multiple-choice questions to assess the LLMs’ understanding of CTI standards, threats, detection strategies, mitigation plans, and best practices. We utilize various authoritative sources and standards within CTI domain such as NIST [29], MITRE [10], GDPR [2] to craft this dataset. Next, we propose two practical CTI tasks that evaluate LLMs’ reasoning and problem-solving skills: mapping Common Vulnerabilities and Exposures (CVE) descriptions [40] to Common Weakness Enumeration (CWE) categories [41] (CTI-RCM) and calculating the Common Vulnerability Scoring System (CVSS) scores (CTI-VSP). Both tasks assess LLMs’ ability to understand and evaluate cyber threats and vulnerabilities. Additionally, we create a task requiring LLMs to analyze publicly available threat reports and attribute the reports to specific threat actors or malware families (CTI-TAA). This task necessitates a thorough understanding of how different cyber threats or malware families have behaved in the past to identify meaningful correlations. Together, these tasks provide a robust framework for assessing LLMs in CTI. Figure 1 provides an overview of our benchmark. We evaluate five state-of-the-art LLMs—three commercial and two open-source—on these tasks. Our results and analysis provide important insights into the LLMs’ capabilities in CTI analysis and highlight future avenues for research. We make the datasets and code publicly available.
@@ 3 CTIBench: A Benchmark for Evaluating LLMs in CTI
We are motivated by the need to create knowledge-intensive tasks to evaluate the cognitive capabilities of LLMs in Cyber Threat Intelligence (CTI) [30]. Our benchmark aims to verify that LLMs have the ability to understand, investigate, and analyze cyber threat reports. To this end, we have designed tasks and datasets that emphasize four fundamental cognitive capabilities of LLMs in CTI: memorization (ability to recall and utilize previously learned information), understanding (ability to comprehend the content and context), problem-solving (ability to apply knowledge and reasoning to address specific challenges), and reasoning (ability to draw logical conclusions and make informed decisions based on available information) [54].
@@ 3.1 CTI-MCQ: Cyber Threat Intelligence Multiple Choice Questions
Data Collection. To generate the CTI-MCQ dataset, we utilize various authoritative sources within the CTI domain. These sources include, among others, CTI frameworks such as NIST [29], the Diamond Model of Intrusion Detection [18], and regulations like GDPR [2]. Additionally, we incorporate CTI sharing standards such as STIX and TAXII [12] to formulate questions on fundamental cybersecurity and CTI knowledge. We leveraged the MITRE ATT&CK framework [46], CWE database [20], and CAPEC [19] to develop questions on attack patterns, threat actors, APT campaigns, detection methods, mitigation strategies, common software vulnerabilities and attack pattern enumeration. Finally, we supplement our dataset by manually collecting and curating questions from publicly available CTI quizzes, ensuring their relevance and accuracy by referencing established CTI resources. While these quizzes may introduce some bias, as they may not fully represent the entire spectrum of CTI knowledge, we ensured our diverse range of sources and manual curation process mitigated this potential limitation.
Generating Questions Using GPT-4. We utilize GPT-4o [6] to prepare the MCQs. Initially, we preprocess the content from the webpage to remove sections inappropriate for MCQ generation (such as page creation or update dates, references, etc.). We then optimize the prompt for creating questions that are challenging enough to test the knowledge of LLMs in cybersecurity. An example prompt is shown in Prompt A. We vary the number of questions we generate based on document length; we create more questions as the length of the document increases. We then randomly sample approximately 3000 questions for manual validation. This approach ensures a variety of questions while retaining quality
Dataset Validation. We manually analyze the quality of MCQs extracted from ChatGPT-4o to ensure the quality of the CTI-MCQ dataset. The human annotators were given access to the source URLs when analyzing the questions. We identified two major issues from LLM-generated questions: some questions had multiple correct options in the answers, and sometimes the LLM-given answer was incorrect. We removed the questions that were unanswerable from the given context or questions that included multiple correct options. We fixed the responses that had incorrect annotations provided by ChatGPT-4o.
@@ 3.2 CTI-RCM: Cyber Threat Intelligence Root Cause Mapping
Data Collection. For this task, we collect data from the National Vulnerability Database (NVD) [11]. The NVD database provides descriptions of past vulnerabilities identified by CVE, along with their associated mappings to Common Weakness Enumeration (CWE) entries. For our study, we specifically focus on vulnerabilities reported in the year 2024 that include associated CWE mappings. We then randomly sample 1,000 vulnerabilities to include in our dataset.
@@ 3.3 CTI-VSP: Cyber Threat Intelligence Vulnerability Severity Prediction
The Vulnerability Severity Prediction task involves predicting the Common Vulnerability Scoring System (CVSS) vector string from a given vulnerability description. The CVSS is a standardized framework used to assess the severity of security vulnerabilities. It is composed of three metric groups: Base, Temporal, and Environmental. The Base metric group, which we focus on in our study, reflects the severity of a vulnerability based on its intrinsic characteristics. These characteristics are constant over time and assume a reasonable worst-case impact across different deployed environments. More details can be found in Appendix B.
@@
Data Collection. We use the same data source as the CTI-RCM for this task. Specifically, we collect 1,000 vulnerability descriptions from 2024 and their associated CVSS v3 strings.
@@ 4.2 Evaluation Metrics
We use accuracy to evaluate the CTI-MCQ and CTT-RCM tasks, as both tasks are equivalent to multi-class classification. For the CTI-VSP task, we compute the mean absolute deviation (MAD) between the CVSS v3.1 scores of the ground truth and the model’s predictions. Although we ask the model to predict a vector string, the CVSS score can be deterministically derived from it. We utilize the Python library cvss [42] to compute the CVSS score (a numerical value in the range of 0-10 that determines the overall severity of a vulnerability) from the predicted string, ensuring that any potential errors from the LLM performing the computation are eliminated. This approach focuses solely on assessing the LLM’s reasoning ability regarding vulnerability.
@@ 4.3 Results Summary
Table 1 presents the performance of various LLMs on our benchmark CTI tasks. The results indicate that GPT-4 outperforms other models across all tasks except for CTI-VSP, where Gemini-1.5 takes the lead. Despite being open-source, LLAMA3-70B performs comparably to Gemini-1.5 and even outperforms it on two tasks, though it struggles with the CTI-VSP task. ChatGPT-3.5 exceeds the performance of LLAMA3-8B but is generally outperformed by the other models across most tasks. LAMA3-8B, being a smaller model, fails to match the performance of larger models on tasks requiring more nuanced understanding and reasoning. However, it performs decently on the CTI-MCQ task.
@@
Model CTI-MCQ (Acc) CTI-RCM (Acc) CTI-VSP (MAD) Correct Plausible
ChatGPT-4 71.0 72.0 1.31 52 86
ChatGPT-3.5 54.1 67.2 1.57 44 62
Gemini-1.5 65.4 66.6 1.09 38 74
LLAMA3-70B 65.7 65.9 1.83 52 80
LLAMA3-8B 61.3 44.7 1.91 28 36
@@ 5.2 CTI-RCM
(c) Accuracy for the CTI-ROM task for 2 different years
@@ 5.4 CTI-TAA
Correct response: We task the LLM to predict the threat actor CHRYSENE given a threat report by replacing the mention of the threat actor and its campaign with [PlaceHolder]. [^2] https://www.welivesecurity.com/en/eset-research/oilrigs-outer-space-juicy-mix-same-ol-rig-new-drill-pipes/ CHRYSENE is an Iranian cyberespionage group active since 2014, targeting Middle Eastern governments and various industries. ChatGPT-4 predicted the threat actor as OilRig, which is an alias of CHRYSENE.
Plausible response: We task the LLM to predict the threat actor MuddyWater given a threat report by replacing the mention of the threat actor and its campaign with [PlaceHolder]. [^3] https://www.deepinstinct.com/blog/darkbeatc2-the-latest-muddywater-attack-framework MuddyWater is a cyber espionage group linked to Iran’s Ministry of Intelligence and Security (MOIS) that targets government and private sectors across the Middle East, Asia, Africa, Europe, and North America. ChatGPT-4 predicted the threat actor as APT35, which is not the alias of MuddyWater but shares some common attack patterns like originating from Iran, targeting the Middle East and North America, and using multi-stage attacks.
@@ 5.5 Compute Cost
Appendix F presents the detailed token counts for each task. GPT-4 generated significantly longer responses across all tasks than the other models.
@@ 7 Limitations
While our evaluation of LLMs for Cyber Threat Intelligence (CTI) tasks provides valuable insights, it is important to acknowledge certain limitations. First, the extensive range of CTI tasks poses a significant challenge, and our study has only selected a limited subset of these tasks to measure the LLM’s capabilities. This selection may not fully represent the comprehensive functionality required for effective CTI operations. We will broaden the scope of evaluated tasks to encompass a wider array of CTI activities, ensuring a more holistic assessment of LLM performance in our future works. Second, our evaluation is restricted to English-language CTI techniques. This limitation overlooks the global nature of cyber threats, which often involve multiple languages. Future studies will consider incorporating multilingual CTI evaluations to better reflect the diverse linguistic landscape of cyber threats and enhance the applicability of LLMs in international contexts.
@@ C.3 CTI-VSP: Cyber Threat Intelligence Vulnerability Severity Prediction
C.4 CTI-TAA: Cyber Threat Intelligence Threat Actor Attribution
@@ Appendix D CTI-TAA Evaluation
The algorithm is shown in Algorithm 1. We make the algorithm included in our released code.
@@
Appendix E CTI-MCQ Samples

### v3

@@ 1 Introduction
Addressing this gap, we propose CTIBench, a novel suite of benchmark tasks and datasets to evaluate LLMs in cyber threat intelligence. To this end, we construct a knowledge evaluation dataset, CTI-MCQ, comprising multiple-choice questions aimed at testing LLMs’ understanding of crucial CTI concepts, including standards, threat identification, detection strategies, mitigation techniques, and best practices. We utilize various authoritative sources and standards within CTI domain such as NIST [30], MITRE [10], GDPR [2] to craft this dataset. In addition, we introduce three practical CTI tasks designed to evaluate LLMs’ reasoning and problem-solving capabilities: (1) CTI-RCM, which involves mapping Common Vulnerabilities and Exposures (CVE) descriptions [41] to Common Weakness Enumeration (CWE) categories [42]; (2) CTI-VSP, which requires calculating Common Vulnerability Scoring System (CVSS) scores [24]; and (3) CTI-ATE, which focuses on extracting MITRE ATT&CK techniques from threat descriptions [15]. These tasks assess the LLMs’ proficiency in understanding and evaluating cyber threats, vulnerabilities, and attack patterns. Furthermore, we propose a more complex task, CTI-TAA, where LLMs are tasked with analyzing publicly available threat reports and attributing them to specific threat actors or malware families. This task necessitates a thorough understanding of how different cyber threats or malware families have behaved in the past to identify meaningful correlations. Together, these tasks provide a robust framework for assessing LLMs in CTI. Figure 1 provides an overview of our benchmark. We evaluate five state-of-the-art LLMs—three commercial and two open-source—on these tasks. Our results and analysis provide important insights into the LLMs’ capabilities in CTI analysis and highlight future avenues for research. We make the datasets and code publicly available.
@@ 3 CTIBench: A Benchmark for Evaluating LLMs in CTI
We are motivated by the need to create knowledge-intensive tasks to evaluate the cognitive capabilities of LLMs in Cyber Threat Intelligence (CTI) [31]. Our benchmark aims to verify that LLMs can understand, investigate, and analyze cyber threat reports. To this end, we have designed tasks and datasets that emphasize four fundamental cognitive capabilities of LLMs in CTI: memorization (ability to recall and utilize previously learned information), understanding (ability to comprehend the content and context), problem-solving (ability to apply knowledge and reasoning to address specific challenges), and reasoning (ability to draw logical conclusions and make informed decisions based on available information) [55].
@@ 3.1 CTI-MCQ: Cyber Threat Intelligence Multiple Choice Questions
Data Collection.
To generate the CTI-MCQ dataset, we utilize various authoritative sources within the CTI domain. These sources include, among others, CTI frameworks such as NIST [30], the Diamond Model of Intrusion Detection [19], and regulations like GDPR [2]. Additionally, we incorporate CTI sharing standards such as STIX and TAXII [12] to formulate questions on fundamental cybersecurity and CTI knowledge. We leveraged the MITRE ATT&CK framework [47], CWE database [21], and CAPEC [20] to develop questions on attack patterns, threat actors, APT campaigns, detection methods, mitigation strategies, common software vulnerabilities and attack pattern enumeration. Finally, we supplement our dataset by manually collecting and curating questions from publicly available CTI quizzes, ensuring their relevance and accuracy by referencing established CTI resources. While these quizzes may introduce some bias, as they may not fully represent the entire spectrum of CTI knowledge, we ensured our diverse range of sources and manual curation process mitigated this potential limitation.
Generating Questions Using GPT-4.
We utilize GPT-4o [6] to prepare the MCQs. Initially, we preprocess the content from the webpage to remove sections inappropriate for MCQ generation (such as page creation or update dates, references, etc.). We then optimize the prompt for creating questions that are challenging enough to test the knowledge of LLMs in cybersecurity. An example prompt is shown in Prompt A. We vary the number of questions we generate based on document length; we create more questions as the length of the document increases. We then randomly sample approximately 3000 questions for manual validation. This approach ensures a variety of questions while retaining quality
Dataset Validation.
We manually analyze the quality of MCQs extracted from ChatGPT-4o to ensure the quality of the CTI-MCQ dataset. The human annotators were given access to the source URLs when analyzing the questions. We identified two major issues from LLM-generated questions: some questions had multiple correct options in the answers, and sometimes the LLM-given answer was incorrect. We removed the questions that were unanswerable from the given context or questions that included multiple correct options. We fixed the responses that had incorrect annotations provided by ChatGPT-4o.
@@ 3.2 CTI-RCM: Cyber Threat Intelligence Root Cause Mapping
Data Collection.
For this task, we collect data from the National Vulnerability Database (NVD) [11]. The NVD database provides descriptions of past vulnerabilities identified by CVE, along with their associated mappings to Common Weakness Enumeration (CWE) entries. For our study, we specifically focus on vulnerabilities reported in the year 2024 that include associated CWE mappings. We then randomly sample 1,000 vulnerabilities to include in our dataset.
@@ 3.3 CTI-VSP: Cyber Threat Intelligence Vulnerability Severity Prediction
The Vulnerability Severity Prediction task involves predicting the Common Vulnerability Scoring System (CVSS) vector string from a given vulnerability description [24]. The CVSS is a standardized framework used to assess the severity of security vulnerabilities. It is composed of three metric groups: Base, Temporal, and Environmental. The Base metric group, which we focus on in our study, reflects the severity of a vulnerability based on its intrinsic characteristics. These characteristics are constant over time and assume a reasonable worst-case impact across different deployed environments. More details can be found in Appendix B.
@@
Data Collection.
We use the same data source as the CTI-RCM for this task. Specifically, we collect 1,000 vulnerability descriptions from 2024 and their associated CVSS v3 strings.
3.4 CTI-ATE: Cyber Threat Intelligence Attack Technique Extraction
The Attack Technique Extraction task involves identifying specific attack patterns from a given description of threat behavior and mapping them to the corresponding MITRE ATT&CK technique IDs [15]. These technique IDs represent distinct adversarial methods used at various stages of an attack lifecycle.
Consider the following example [^2] https://attack.mitre.org/software/S0163/:
Janicab is an OS X trojan that exploited a valid Apple Developer ID to deceive users during installation. It captured both audio and screenshots, which were then transmitted to a remote command and control (C2) server. For persistence, Janicab employed a cron job on affected Mac devices. The use of a legitimate Apple Developer ID enabled the trojan to bypass security restrictions by signing the malicious code.
From this description, we can identify the following relevant attack technique IDs: (i) T1123 – Audio Capture, (ii) T1053 – Scheduled Task, (iii) T1113 – Screen Capture, and (iv) T1553 – Subvert Trust Controls.
This task is valuable for CTI practitioners, as accurately mapping observed behaviors to the corresponding MITRE ATT&CK techniques is essential for designing effective mitigation strategies and deploying targeted security measures. By linking specific behaviors to their respective technique IDs, security teams can better understand the adversary’s tactics, techniques, and procedures (TTPs), enabling them to take proactive actions to disrupt or mitigate ongoing threats.
Data Collection.
For this task, we curated a dataset using information from the MITRE ATT&CK framework [15], which provides comprehensive descriptions of various malware and their associated adversarial behaviors, each categorized by a unique attack technique ID based on open-source threat reports. Our dataset includes 30 instances of malware reported in 2024, alongside their corresponding attack technique IDs—information that extends beyond the knowledge cutoff of the LLMs under evaluation. We also supplemented the dataset with 30 malware instances reported before 2024. For this task, we focused solely on techniques (excluding sub-techniques). In total, the dataset comprises 397 unique attack techniques.
@@ 4.2 Evaluation Metrics
We use accuracy to evaluate the CTI-MCQ and CTT-RCM tasks, as both tasks are equivalent to multi-class classification. For the CTI-VSP task, we compute the mean absolute deviation (MAD) between the CVSS v3.1 scores of the ground truth and the model’s predictions. Although we ask the model to predict a vector string, the CVSS score can be deterministically derived from it. We utilize the Python library cvss [43] to compute the CVSS score (a numerical value in the range of 0-10 that determines the overall severity of a vulnerability) from the predicted string, ensuring that any potential errors from the LLM performing the computation are eliminated. This approach focuses solely on assessing the LLM’s reasoning ability regarding vulnerability. We adopt the Micro-F1 score as the evaluation metric for the CTI-ATE task. Given that this task requires accurately extracting the relevant attack techniques from the provided text and mapping them to their corresponding MITRE ATT&CK technique IDs, the Micro-F1 score is suitable for capturing both precision and recall in a multi-label classification setting.
@@ 4.3 Results Summary
Table 1 presents the performance of various LLMs on our benchmark CTI tasks. The results indicate that GPT-4 outperforms other models across all tasks except for CTI-VSP, where Gemini-1.5 takes the lead. Despite being open-source, LLAMA3-70B performs comparably to Gemini-1.5 and even outperforms it on three tasks, though it struggles with the CTI-VSP task. ChatGPT-3.5 exceeds the performance of LLAMA3-8B but is generally outperformed by the other models across most tasks. LAMA3-8B, being a smaller model, fails to match the performance of larger models on tasks requiring more nuanced understanding and reasoning. However, it performs decently on the CTI-MCQ task.
@@
Model CTI-MCQ (Acc) CTI-RCM (Acc) CTI-VSP (MAD) CTI-ATE (Macro-F1) Correct Plausible
ChatGPT-4 71.0 72.0 1.31 0.6388 52 86
ChatGPT-3.5 54.1 67.2 1.57 0.3108 44 62
Gemini-1.5 65.4 66.6 1.09 0.4612 38 74
LLAMA3-70B 65.7 65.9 1.83 0.4720 52 80
LLAMA3-8B 61.3 44.7 1.91 0.1562 28 36
@@ 5.1 CTI-MCQ
While Table 1 presents the results for the MCQ task without an explicit reasoning step, we also performed an additional evaluation incorporating an explicit reasoning prompt. The detailed results of this evaluation are provided in Appendix E. However, this approach did not consistently lead to performance improvements. We hypothesize that this is because the MCQ task primarily relies on memorization rather than reasoning.
@@ 5.2 CTI-RCM
(c) Accuracy for the CTI-RCM task for 2 different years
@@ 5.3 CTI-VSP
5.4 CTI-ATE
Model Before (F1) After (F1)
ChatGPT-4 0.6542 0.6208
ChatGPT-3.5 0.3420 0.3333
Gemini-1.5 0.4360 0.5263
LLAMA3-70B 0.4934 0.4297
LLAMA3-8B 0.1813 0.1366
Table 2: Performance comparison on the CTI-ATE task, evaluated on samples before and after the models’ knowledge cutoff dates.
The results in Table 1 demonstrate that ChatGPT-4 significantly outperforms other models on the CTI-ATE task, underscoring the complexity of the task. To further investigate model performance, we evaluated the models on samples collected before and after their respective knowledge cutoff dates. These results are presented in Table 2. Most models exhibit slightly better performance on samples collected before their knowledge cutoff, except for Gemini, which performs better on post-cutoff samples. The performance differences are generally insignificant, suggesting that this can serve as a reasonable indicator of how LLMs will fare on future threat reports.
@@ 5.5 CTI-TAA
Correct response:
We task the LLM to predict the threat actor CHRYSENE given a threat report by replacing the mention of the threat actor and its campaign with [PlaceHolder]. [^3] https://www.welivesecurity.com/en/eset-research/oilrigs-outer-space-juicy-mix-same-ol-rig-new-drill-pipes/ CHRYSENE is an Iranian cyberespionage group active since 2014, targeting Middle Eastern governments and various industries. ChatGPT-4 predicted the threat actor as OilRig, which is an alias of CHRYSENE.
Plausible response:
We task the LLM to predict the threat actor MuddyWater given a threat report by replacing the mention of the threat actor and its campaign with [PlaceHolder]. [^4] https://www.deepinstinct.com/blog/darkbeatc2-the-latest-muddywater-attack-framework MuddyWater is a cyber espionage group linked to Iran’s Ministry of Intelligence and Security (MOIS) that targets government and private sectors across the Middle East, Asia, Africa, Europe, and North America. ChatGPT-4 predicted the threat actor as APT35, which is not the alias of MuddyWater but shares some common attack patterns like originating from Iran, targeting the Middle East and North America, and using multi-stage attacks.
@@
Model Before (Acc) After (Acc)
Correct Plausible Correct Plausible
ChatGPT-4 58.06 90.32 42.10 78.95
ChatGPT-3.5 50.00 75.00 43.48 60.87
Gemini-1.5 34.61 76.92 41.66 70.83
LLAMA3-70B 58.06 83.87 42.10 73.68
LLAMA3-8B 25.00 25.00 28.94 39.47
Table 3: Performance comparison of models before and after knowledge cutoff dates for the CTI-TAA task.
Impact of Knowledge Cutoff:
We evaluate the LLMs on the CTI-TAA task using datasets from both before and after their knowledge cutoff dates. The results are displayed in the table below. With the exception of LLAMA3-8B, all other LLMs demonstrated better performance on datasets available during their training period, suggesting that memorization plays a role in improving performance to some extent in this task.
@@ 5.6 Compute Cost
Appendix G presents the detailed token counts for each task. GPT-4 generated significantly longer responses across all tasks than the other models.
@@ 7 Limitations
While our evaluation of large language models (LLMs) for Cyber Threat Intelligence (CTI) tasks provides valuable insights, it is important to recognize certain limitations. First, the extensive scope of CTI activities presents a significant challenge, and our study has focused on a limited subset of tasks to assess the capabilities of LLMs. This selection may not fully capture the breadth of functionality required for comprehensive CTI operations. In future work, we plan to expand the range of evaluated tasks to encompass a broader spectrum of CTI activities, thereby ensuring a more holistic assessment of LLM performance in this domain.
Second, our evaluation is restricted to English-language CTI techniques. This limitation neglects the global nature of cyber threats, which frequently span multiple languages and regions. To address this, future studies will incorporate multilingual CTI evaluations, better reflecting the diverse linguistic landscape of cyber threats and improving the applicability of LLMs in international cybersecurity contexts.
Additionally, there is a genuine risk of the malicious use of LLMs to exploit CTI knowledge for harmful purposes. For instance, if misused, LLMs could generate and disseminate compelling but false threat intelligence reports. Such reports might mislead decision-makers, result in the misallocation of resources, or prompt inappropriate responses. Benchmarking the potential for such misuse remains an open area for future research.
@@ C.3 CTI-VSP: Cyber Threat Intelligence Vulnerability Severity Prediction
C.4 CTI-ATE: Cyber Threat Intelligence Attack Technique Extraction
Prompt used for LLM Evaluation
We provide the list of all Enterprise attack techniques in the prompt for the CTI-ATE task:
You are a cybersecurity expert specializing in cyber threat intelligence. Extract all MITRE Enterprise attack patterns from the following text and map them to their corresponding MITRE technique IDs. Provide reasoning for each identification. Ensure the final line contains only the IDs for the main techniques, separated by commas, excluding any subtechnique IDs. MITRE Enterprise IDs are given below as reference. Text: List of All MITRE Enterprise technique IDs
C.5 CTI-TAA: Cyber Threat Intelligence Threat Actor Attribution
@@ Appendix D CTI-TAA Evaluation
The algorithm is shown in Algorithm 1. We include the algorithm in our released code.
@@
Appendix E CTI-MCQ with reasoning
Model Acc Without Reasoning Acc With Reasoning
ChatGPT-4 71.00 (105) 71.84 (618)
ChatGPT-3.5 54.08 (10) 59.16 (227)
Gemini-1.5 65.44 (5) 65.68 (360)
LLAMA3-70B 65.72 (1) 65.48 (230)
LLAMA3-8B 61.32 (5) 55.80 (269)
Table 4: Comparison of model accuracy with and without reasoning. Numbers in parentheses indicate the approximate number of tokens generated (in thousand tokens)
We re-evaluated the LLMs on the MCQ task by modifying the prompt to explicitly instruct the models to reason through the problem before providing an answer. Table 4 presents the results. While explicit reasoning prompts substantially increased computational costs, they generally do not translate into improved performance. As shown in the table, only GPT-3.5 exhibited a notable performance boost, while for LLAMA3-8B, the reasoning prompt led to a decline in performance.
Appendix F CTI-MCQ Samples